OpenStudy (anonymous):

solve each system using substitution. check your answer. y=8-x 7=2-y

OpenStudy (solomonzelman):

maybe you meant: \(\normalsize\color{royalblue}{ \rm y=8-x }\) \(\normalsize\color{royalblue}{ \rm x=2-y }\) ???

OpenStudy (anonymous):

no thats how it is

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm y=8-x }\) \(\normalsize\color{royalblue}{ \rm 7=2-y }\) solve for y, in equation 2. Substitute your y solution into the first equation. solve for x in the (new formed) first equation

OpenStudy (anonymous):

Solve for one of the variables first. Im going to solve for y in the bottom first. Using the subtraction and multiplication properties of equality. You'll get that y= -5. Substitute that into the first one and solve for x.